# Cristian Sestito  Shady Agwa  and Themis Prodromakis

### v1

@@ Abstract
In order to follow the ever-growing computational complexity and data intensity of state-of-the-art AI models, new computing paradigms are being proposed. These paradigms aim at achieving high energy efficiency, by mitigating the Von Neumann bottleneck that relates to the energy cost of moving data between the processing cores and the memory. Convolutional Neural Networks (CNNs) are particularly susceptible to this bottleneck, given the massive data they have to manage. Systolic Arrays (SAs) are promising architectures to mitigate the data transmission cost, thanks to high data utilization carried out by an array of Processing Elements (PEs). These PEs continuously exchange and process data locally based on specific dataflows (like weight stationary and row stationary), in turn reducing the number of memory accesses to the main memory. The hardware specialization of SAs can meet different workloads, ranging from matrix multiplications to multi-dimensional convolutions. In this paper, we propose TrIM: a novel dataflow for SAs based on a Triangular Input Movement and compatible with CNN computing. When compared to state-of-the-art SA dataflows, like weight stationary and row stationary, the high data utilization offered by TrIM guarantees 10$\sim$ 10$\times$ 10 × less memory access. Furthermore, considering that PEs continuously overlap multiplications and accumulations, TrIM achieves high throughput (up to 81.8% higher than row stationary), other than requiring a limited number of registers (up to 15.6×15.6$\times$15.6 × fewer registers than row stationary).
@@ I Introduction
To address the drawbacks of previous art, we propose TrIM, a new dataflow that exploits a triangular input movement to maximize input utilization. Advantageously, this results in less memory accesses without the need to introduce data redundancy. A generic SA dealing with TrIM consists of $K\times K$ PEs, where weights are kept stationary and psums are propagated vertically and finally accumulated by an extra adder tree. The TrIM-based SA is compatible with the computation requirements of Convolutional Layers (CLs). When compared to WS, TrIM exhibits one order of magnitude less memory accesses. Moreover, TrIM ensures high performance efficiency since each PE works at the peak throughput of 2 OPs/cycle. Finally, the micro-architecture of PEs is fairly simple, translating into a limited number of registers, up to 15.6×15.6$\times$15.6 × lower than RS with a kernel size of 7.
@@
Conv-to-GeMM Convolution to General Matrix Multiplication Conversion
@@ III Triangular Input Movement Systolic Array
The TrIM-based SA consists of $K\times K$ PEs, where weights are retained at the PE level, while inputs and psums are moved across the array to maximize data reuse. Before any computation, weights are read from the memory and provided to the PEs placed at the top side of the array; then, these are moved from top to down to allow the other weights to be accommodated. Inputs supply the PEs through a novel data movement that can be summarized in three steps: (i) first, inputs are fetched from the main memory and delivered to the PEs; (ii) then, such inputs move from right to left, until they reach the left edge of the array; (iii) finally, they move diagonally towards the upper PEs. According to the width of the ifmap under processing, the leftmost PEs may be connected to Shift Register Buffers (SRBs) with depth $W_{I}-K-1$, which ensure the correct execution of the diagonal movement over time. Steps (i) to (iii) compose a right-triangular shape, hence the name Triangular Input Movement (TrIM). Fig. 4 schematizes a generic architecture coping with the proposed dataflow. Each PE is labelled as $PE_{i,j}$, with $0\leq i,j<K$, and consists of a MAC unit, four registers, and two multiplexers to establish whether the current input is reused from a different PE or not. If reused, inputs can move from right to left (thus from each $PE_{i,j+1}$ to $PE_{i,j}$), or they can be forwarded diagonally. In the latter case, they are first provided to the SRBs from each $PE_{i,0}$, then shifted along such buffers, and finally the inputs stored into the last $K$ registers are supplied to the top $PE_{i-1,j}$ elements. Psums are accumulated vertically, thus from each $PE_{i,j}$ to each $PE_{i+1,j}$. Eventually, an adder tree accumulates the psums coming from the the bottom $PE_{K-1,j}$ elements. For very small ifmaps, with $W_{I}\leq 2K$, the number of registers may be lower than $K$, thus the diagonal connections may also interest some of the PEs.
@@ IV-A Modelling WS and RS Dataflows
The number of memory accesses (MA) of the WS-based SA is dictated by Conv-to-GeMM and reported in equation (2):
@@
This conversion results in data redundancy and higher memory capacity, since ifmap’s activations are arranged as $H_{O}\times W_{O}$ rows and $K^{2}$ columns. On the contrary, since no Conv-to-GeMM is required by the RS-based SA, the MA of the main memory are directly related to the ifmap’s sizes ($H_{I}\times W_{I}$). However, the RS-based SA needs memory blocks at the PE level, usually implemented as SRAM-based scratch pads, which severely affect the energy footprint. Based on the Eyeriss implementation [22], the scratch pads energy is estimated to be from $\alpha=\sim 11\times$ to 16.3$\sim$ 16.3$\times$ 16.3 × higher than the memory energy. Therefore, the number of MA is summarized by equation (3):
@@ IV-B Modelling TrIM Dataflow
$Reg_{TrIM}=3\times K^{2}+(K-1)\times W_{I}+2$ (18)
@@ V-A Memory Accesses
In the case of the RS dataflow, the total number of MA is given by two contributions: the main memory and the memory blocks for data circulation at the PE level. From the main memory, inputs are read once, therefore the MA corresponds to the number of activations of each ifmap. This contribution is reported in Fig. 6(a)-(c). TrIM’s MA are practically close to the RS’ main memory accesses. In fact, the overhead introduced by TrIM and previously reported in equation (14) is minimal. For example, when $K=3$ and $I=256$, TrIM uses only 1.5% more MA than RS. Fig. 7 illustrates the overhead required by RS and dictated by the use of memory blocks at the PE level, usually in the form of scratch pads [22] . This contribution is from 11$\sim$ 11$\times$ 11 × to 16.3$\sim$ 16.3$\times$ 16.3 × higher than the memory accesses required to read data.
@@ V-C Registers
Fig. 9 shows the total number of registers for each dataflow. Fig. 9(a) refers to $K=3$, Fig. 9(b) refers to $K=5$, while Fig. 9(c) refers to $K=7$. The RS-based SA requires the highest number of registers, considering that memory blocks for inputs and weights must be accommodated into $K\times H_{O}$ PEs. For example, when $K=3$ and $I=256$, the registers of RS are 10$\sim$ 10$\times$ 10 × more than those required by TrIM. Since the majority of RS’ registers are memory blocks, the degradation in terms of energy is even worse.
@@ V-D Discussion
The design space analysis presented in this Section allow us to highlight several insights:
@@ VI Conclusion
This paper introduces TrIM, an innovative dataflow for SAs that is compatible with CNN computing. The dataflow has two levels of granularity: at the PE level, weights are kept stationary; high inputs utilization is guaranteed at the SA level, based on a triangular movement. To achieve this, $K\times K$ PEs are interconnected with each other in different directions, and local shift registers are put at the left edge of the array to assist the triangular movement of inputs. We built an analytical model to characterize the array before the physical implementation phase. Memory accesses, throughput and number of registers are subjected to in-depth design space exploration. When compared to WS dataflow, TrIM requires one order of magnitude less memory accesses considering that no data redundancy is exploited. Furthermore, TrIM requires less memory access than RS since no micro-architectural memory blocks assist data circulation. The simpler PEs in TrIM also result in 15.6×15.6$\times$15.6 × fewer registers than RS. Finally, thanks to the overlap between multiplications and accumulations, TrIM’s real throughput achieves the peak throughput of 2 OPs/cycle/PE, overcoming RS by 81.8%. In the second part, we present a complete architecture dealing with TrIM and compatible with multi-dimensional convolutions for CNNs. As a case study, the architecture is characterized considering the VGG-16 CNN[29], and compared with state-of-the-art competitors.

### v2

# TrIM, Triangular Input Movement Systolic Array for Convolutional Neural Networks:

@@ Abstract
In order to follow the ever-growing computational complexity and data intensity of state-of-the-art AI models, new computing paradigms are being proposed. These paradigms aim at achieving high energy efficiency, by mitigating the Von Neumann bottleneck that relates to the energy cost of moving data between the processing cores and the memory. Convolutional Neural Networks (CNNs) are susceptible to this bottleneck, given the massive data they have to manage. Systolic Arrays (SAs) are promising architectures to mitigate the data transmission cost, thanks to high data utilization of Processing Elements (PEs). These PEs continuously exchange and process data locally based on specific dataflows (like weight stationary and row stationary), in turn reducing the number of memory accesses to the main memory. In SAs, convolutions are managed either as matrix multiplications or exploiting the raster-order scan of sliding windows. However, data redundancy is a primary concern affecting area, power and energy. In this paper, we propose TrIM: a novel dataflow for SAs based on a Triangular Input Movement and compatible with CNN computing. TrIM maximizes the local input utilization, minimizes the weight data movement and solves the data redundancy problem. Furthermore, TrIM does not incur the significant on-chip memory penalty introduced by the row stationary dataflow. When compared to state-of-the-art SA dataflows the high data utilization offered by TrIM guarantees 10$\sim$ 10$\times$ 10 × less memory access. Furthermore, considering that PEs continuously overlap multiplications and accumulations, TrIM achieves high throughput (up to 81.8% higher than row stationary), other than requiring a limited number of registers (up to 15.6×15.6$\times$15.6 × fewer registers than row stationary).
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
To address the drawbacks of previous art, we propose TrIM, a new dataflow that exploits a triangular input movement to maximize input utilization. A generic SA dealing with TrIM consists of $K\times K$ PEs, where weights are kept stationary and psums are propagated vertically and finally accumulated by an extra adder tree. The TrIM-based SA is compatible with the computation requirements of Convolutional Layers (CLs). While we took inspiration from the WS dataflow to keep weights fixed throughout the computations, we considered the convolution-oriented abstraction of the RS dataflow to overcome the Conv-to-GeMM transformation. However, TrIM sits completely apart from the above referenced dataflows because it solves the data redundancy problem: differently from the baseline WS, TrIM does not require Conv-to-GeMM conversion. Also, TrIM is not penalized by the significant energy requirements associated with the on-chip memory blocks unlike the RS dataflow. Furthermore, TrIM avoids the data redundancy at the array level presented by RS. When compared to WS, TrIM exhibits one order of magnitude less memory accesses. Moreover, TrIM ensures high performance efficiency since each PE works at the peak throughput of 2 OPs/cycle, which results in 81.8% improvement of RS. Finally, the micro-architecture of PEs is fairly simple, translating into a limited number of registers, up to 15.6×15.6$\times$15.6 × lower than RS with a kernel size of 7.
@@
Conv-to-GeMM Convolution to GeMM Conversion
@@ III Triangular Input Movement Systolic Array
III-A The Dataflow
The TrIM-based SA consists of $K\times K$ PEs, where weights are retained at the PE level, while inputs and psums are moved across the array to maximize data reuse. Before any computation, weights are read from the memory and provided to the PEs placed at the top side of the array; then, these are moved from top to down to allow the other weights to be accommodated. Inputs supply the PEs through a novel data movement that can be summarized in three steps: (i) first, inputs are fetched from the main memory and delivered to the PEs; (ii) then, such inputs move from right to left, until they reach the left edge of the array; (iii) finally, they move diagonally towards the upper PEs. According to the width of the ifmap under processing, the leftmost PEs may be connected to Shift Register Buffers (SRBs) with depth $W_{I}-K-1$, which ensure the correct execution of the diagonal movement over time. Steps (i) to (iii) compose a right-triangular shape, hence the name Triangular Input Movement (TrIM). Fig. 4 schematizes a generic architecture coping with the proposed dataflow. Each PE is labelled as $PE_{i,j}$, with $0\leq i,j<K$, and consists of a MAC unit, four registers, and two multiplexers to establish whether the current input is reused from a different PE or not. If reused, inputs can move from right to left (thus from each $PE_{i,j+1}$ to $PE_{i,j}$), or they can be forwarded diagonally. In the latter case, they are first provided to the SRBs from each $PE_{i,0}$, then shifted along such buffers, and finally the inputs stored into the last $K$ registers are supplied to the top $PE_{i-1,j}$ elements. In general, when the generic input is grabbed from the memory, the data $I_{ext}$ is supplied to the PE; when the input stored in the right-adjacent PEs is moved into the current PE, thus the data $I_{R}$ is the target input; finally, when a diagonal movement of the input (either from a bottom PE or from a bottom SRB) has to feed the PE, thus the data $I_{D}$ is provided. Psums are accumulated vertically, thus from each $PE_{i,j}$ to each $PE_{i+1,j}$. Eventually, an adder tree accumulates the psums coming from the the bottom $PE_{K-1,j}$ elements. For very small ifmaps, with $W_{I}\leq 2K$, the number of registers may be lower than $K$, thus the diagonal connections may also interest some of the PEs.
@@ III-A The Dataflow
III-B Differences with WS and RS Dataflows
To better highlight what makes TrIM different from WS-based and RS-based SAs, we list below the key highlights in terms of different features.
  • Data redundancy: WS has data redundancy at the main memory level which adds a significant energy overhead. RS has data redundancy at the array level through memory blocks adding still significant energy overhead. While RS moves the redundancy penalty from the off-chip to the on-chip level, TrIM overcomes the input redundancy problem and moreover maximizes the data utilization by recycling the same input as much as possible.
  • Weights reuse: In WS and TrIM, weights are preliminary fetched from the memory and then fixed at the PE level for the entire duration of the convolution (weight-stationary). In RS, weights are broadcast horizontally to a row of PEs. Each PE receives a set of $K$ weights that in turn are circulated through a scratch pad. WS and TrIM do not require any scratch pad at the PE level, thus making their micro-architecture simpler with a lower energy footprint.
  • Inputs reuse: In WS, inputs are grabbed from the memory and forwarded horizontally to PEs through FIFOs for proper synchronization. In RS, inputs are broadcast diagonally to multiple PEs. Each PE receives a set of $K$ inputs, circulated through scratch pads as long as required. TrIM introduces innovation on the input movement. After being retrieved from the memory and supplied vertically to PEs, inputs are reused through either horizontal and diagonal movements between PEs. The diagonal movements may be assisted by SRBs when the ifmap is fairly larger than the kernel size. Differently from Eyeriss, the diagonal links in TrIM are between PEs or between SRBs and PEs. No broadcasting is thus exploited.
  • Psums propagation: In all the dataflows (WS, RS, TrIM), psums move vertically from top to bottom. However, WS requires FIFOs at the bottom for proper output synchronization, while RS makes the PEs’ micro-architecture complex: the adder has both a feedback loop (to behave as an accumulator during the processing of a row of $K$ inputs and weights) and a vertical connection for psums accumulation between PEs.
  • Array sizes: WS and TrIM make use of $K\times K$ PEs. On the contrary, RS is a rectangular array with sizes $K\times H_{O}$, thus exhibiting more limitations regarding the design scalability and area efficiency. In addition, the presence of scratch pads at the PE level makes the RS even bigger.
@@ IV-A Modelling WS and RS Dataflows
In what follows, memory accesses are to be considered as generalized accesses, without specific reference on the hierarchy level. The number of memory accesses (MA) of the WS-based SA is dictated by Conv-to-GeMM and reported in equation (2):
@@
This conversion results in data redundancy and higher memory capacity, since ifmap’s activations are arranged as $H_{O}\times W_{O}$ rows and $K^{2}$ columns. On the contrary, since no Conv-to-GeMM is required by the RS-based SA, the MA of the main memory are directly related to the ifmap’s sizes ($H_{I}\times W_{I}$). However, the RS-based SA needs memory blocks at the PE level, usually implemented as SRAM-based scratch pads, which severely affect the energy footprint. Based on the Eyeriss implementation [22], the scratch pads energy is estimated to be from $\alpha=\sim 12.9\times$ to 16.5$\sim$ 16.5$\times$ 16.5 × higher than the main memory energy. Therefore, the number of MA is summarized by equation (3):
@@
The $\alpha$ factor is retrieved from [22], considering the total energy dissipated by the scratch pads on the best and worst case scenarios of the VGG-16 CNN [27] and normalized on the total off-chip memory energy. Regarding the off-chip memory energy, we have considered [28] to determine the energy per single access.
@@ IV-B Modelling TrIM Dataflow
$Reg_{TrIM}=4\times K^{2}+(K-1)\times(W_{I}-K-1)+1$ (18)
The component $4\times K^{2}$ relates to the registers available at the PE level (4 for each PE). The second block $(K-1)\times(W_{I}-K-1)$ refers to the registers that model the SRBs. Finally the extra register equips the adder tree.
@@ V-A Memory Accesses
In the case of the RS dataflow, the total number of MA is given by two contributions: the main memory and the memory blocks for data circulation at the PE level. From the main memory, inputs are read once, therefore the MA corresponds to the number of activations of each ifmap. This contribution is reported in Fig. 6(a)-(c). TrIM’s MA are practically close to the RS’ main memory accesses. In fact, the overhead introduced by TrIM and previously reported in equation (14) is minimal. For example, when $K=3$ and $I=256$, TrIM uses only 1.5% more MA than RS. Fig. 7 illustrates the overhead required by RS and dictated by the use of memory blocks at the PE level, usually in the form of scratch pads [22] . This contribution is from 11$\sim$ 11$\times$ 11 × to 16.3$\sim$ 16.3$\times$ 16.3 × higher than the memory accesses required to read data. For example, the factor $\times1$ comes from 12.9/1.169, where $\alpha=12.9$ (Equation 3), while 1.169 relates the normalized memory accesses in TrIM (the 0.169 overhead comes from the analytical model, Equations 13 and 14).
@@ V-C Registers
Fig. 9 shows the total number of registers for each dataflow. Fig. 9(a) refers to $K=3$, Fig. 9(b) refers to $K=5$, while Fig. 9(c) refers to $K=7$. As explained in Section IV, the number of registers provides the local storing capacity of the different dataflows. Thus, simple registers, FIFOs, scratch pads and shift registers are all included in the count. The RS-based SA requires the highest number of registers, considering that memory blocks for inputs and weights must be accommodated into $K\times H_{O}$ PEs. For example, when $K=3$ and $I=256$, the registers of RS are 10$\sim$ 10$\times$ 10 × more than those required by TrIM. Since the majority of RS’ registers are memory blocks, the degradation in terms of energy is even worse.
@@ V-D Discussion
The design space analysis presented in this Section allows us to highlight several insights:
@@ VI Conclusion
This paper introduces TrIM, an innovative dataflow for SAs that is compatible with CNN computing. The dataflow has two levels of granularity: at the PE level, weights are kept stationary; high inputs utilization is guaranteed at the SA level, based on a triangular movement. To achieve this, $K\times K$ PEs are interconnected with each other in different directions, and local shift registers are put at the left edge of the array to assist the triangular movement of inputs. We built an analytical model to characterize the array before the physical implementation phase. Memory accesses, throughput and number of registers are subjected to in-depth design space exploration. When compared to WS dataflow, TrIM requires one order of magnitude less memory accesses considering that no data redundancy is exploited. Furthermore, TrIM requires less memory access than RS since no micro-architectural memory blocks assist data circulation. The simpler PEs in TrIM also result in 15.6×15.6$\times$15.6 × fewer registers than RS. Finally, thanks to the overlap between multiplications and accumulations, TrIM’s real throughput achieves the peak throughput of 2 OPs/cycle/PE, overcoming RS by 81.8%. In[31], we present a complete architecture dealing with TrIM and compatible with multi-dimensional convolutions for CNNs. The architecture is made reconfigurable to support different ifmap sizes and kernel sizes. Implemented onto a FPGA, the design is characterized in terms of resource utilization, performance, energy efficiency. A thorough analysis of the off-chip and on-chip memory access count is reported considering two CNN models. Comparisons with different dataflows and other FPGA implementations are reported, to highlight the advantages of the proposed TrIM-based systolic array.